# Probabilistic Predictions of Process-Induced Deformation in Carbon/Epoxy Composites Using a Deep Operator Network

### v1

@@ 1 Introduction
Along with advances in physics-based modeling, rapid progress in machine learning (ML) has transformed manufacturing research by enabling efficient data-driven modeling, prediction, and optimization [behbahani2022machine, ademujimi2017review, park2025machine, kiyani2023designing]. Deep learning, in particular, has become a central tool across science and engineering. A growing body of work focuses on operator learning, which aims to approximate mappings between input and output functions, a setting that naturally arises in many scientific and engineering applications. For PID in composite structures, a range of ML architectures has been explored. Lin et al. [lin2021use] used neural networks, support vector regression, and $k$-nearest neighbors models to rapidly estimate differences in degree of cure from process parameters, while Fan et al. [fan2023deep] employed convolutional neural networks to predict PID fields for laminates with varying stacking sequences. Theory-guided and simulation-informed approaches have also been proposed, including theory-guided ML with Gaussian process regression for process optimization [schoenholz2024accelerated], random forests for optimal cure design in co-cured composite structures [2022_Lavaggi], and simulation frameworks that capture competing mechanisms of residual stress evolution [18_Limaye2024]. More recently, Liu et al. [liu2024rapid] combined viscoelastic finite element analysis, feature selection (Random Forest, LASSO), and artificial neural networks for rapid PID prediction, while Zhang et al. [zhang2025deep] introduced deep learning models tailored to complex thermo–chemo–mechanical interactions.
@@
Building on these developments, we extend Feature-wise Linear Modulation (FiLM) [perez2018film] to operator learning. FiLM was originally introduced for visual reasoning and has since been integrated into architectures such as graph neural networks [brockschmidt2020gnn]. In this work, a FiLM-augmented DeepONet serves as a surrogate model to reconstruct the full histories of degree of cure, viscosity, and deformation, even though experimentally only the final deformation at the end of the cure cycle is observed. To reconcile the mismatch between the time discretization used in simulations and that available in experiments, we employ a transfer learning strategy [pan2020transfer] where the DeepONet is first trained on high-fidelity simulation data and subsequently fine-tuned on experimental measurements via a loss function that enforces agreement between the predicted and measured final deformation.
@@
While the FiLM DeepONet framework provides accurate predictions for the evolution of degree of cure, viscosity, and deformation, performing reliable uncertainty quantification (UQ) for operator learning remains substantially more challenging. The associated inference problem is typically a very high-dimensional Bayesian inverse problem, which renders gradient-based sampling methods prohibitively expensive. Classical Bayesian approaches such as Hamiltonian Monte Carlo (HMC) [neal2012bayesian] and the Laplace approximation [mackay1992practical] can, in principle, deliver high-quality uncertainty estimates, but they scale poorly, even when diagonal approximations of the Hessian are employed. As a practical alternative, Lakshminarayanan et al. [lakshminarayanan2017simple] proposed deep ensembles trained with proper scoring rules, which can yield well-calibrated predictive distributions with only minor modifications to standard training pipelines. However, these ensemble methods have been criticized for lacking a clear Bayesian interpretation [wilson2022evaluating, reich2013nonparametric, pearce2018uncertainty]. To help bridge this gap, Pearce et al. [pearce2020uncertainty] introduced an ensemble scheme based on approximate Bayesian inference, in which model parameters are regularized around samples from a prior-like distribution, leading to more interpretable uncertainty estimates.
@@
Another promising direction is ensemble Kalman inversion (EKI), a gradient-free ensemble method that has been successfully combined with UQ in scientific machine learning [gao2025scalable, pensoneault2025uncertainty]. Owing to its favorable scaling and rapid convergence, EKI is particularly well suited to high-dimensional operator-learning problems. In this work, we therefore employ EKI to train DeepONet while simultaneously obtaining uncertainty estimates. To further reduce the effective parameter dimension and enhance computational efficiency, we replace conventional multilayer perceptrons (MLPs) with Chebyshev-enhanced Kolmogorov–Arnold Networks (cKANs) within the DeepONet architecture. The cKANs used here follow a modified DeepOKAN design, which incorporates Chebyshev polynomials into Kolmogorov–Arnold Networks (KANs) [liu2024kan, shukla2024comprehensive, faroughi2025scientific] in a standard DeepONet framework. Inspired by Kolmogorov networks [sprecher2002space, koppen2002training], KANs dynamically adapt their activation patterns to the input data, enabling more efficient forward evaluations and, consequently, faster generation of ensemble samples. The model developed as part of this study, thus couples operator learning with transfer learning by training a DeepONet surrogate on simulation data and refining it with experimental measurements to recover full deformation fields from limited observations, while EKI-based training simultaneously yields uncertainty estimates, providing experimental predictions that are both accurate and uncertainty-aware.
@@ 2.1 Experimental Approach
The laminate layup was carried out with particular care to ensure consistent ply orientation across all specimens, thereby reducing variability associated with fiber misalignment. Samples fabricated from the selected thermoset prepreg were cured under constant consolidation pressure using one baseline isothermal cycle and two optimized non-isothermal cycles aimed at mitigating PID. The temperature histories of the specimens were recorded in real time and subsequently used to predict the deformation history using a FiLM-enhanced DeepONet. .
@@ 2.2 Materials and Methods
White et al. [25_White1993] 1993
AS4 CF/3501-6 Prepreg
(42% resin mass)
@ 20 °C/min to 350 °C
182.87 (prepreg)
435 (comp. for resin)
NA
Kim et al. [21_Kim2002] 2002
AS4 CF/3501-6 Prepreg
(37.4% resin mass)
@ 20 °C/min to 350 °C 433.7 (comp. for resin) NA
Kim et al. [21_Kim2002] 2002
AS4 CF/3501-6 Prepreg
(37.4% resin mass)
@ 2 °C/min to 350 °C 456.7 (comp. for resin) NA
Chern et al. [23_Chern2002] 2002 3501-6 Resin @ 20 °C/min $508\pm 19$
2% cure
(manufacturer’s note)
Hargis et al. [20_Hargis2006] 2006 3501-6 Resin @ 20 °C/min $382.5\pm 20$
Significant
pre-reaction reported
@@
Based on the values reported in Table 1 and the method for determining the initial degree of cure from Equation 1, $\alpha$ was estimated to lie within the range of 17.68%–31.21%, considering the standard deviations of the measurements. Using the reported average values, the degree of cure was calculated to be 24.7%. Accordingly, an initial condition of $\alpha=0.3$ (30%) was assumed for the simulations. This approximation also accounts for further curing that may have occurred in the several months between the manufacturing of the prepreg and use for experiments.
@@ 3 Model and Simulations
A comprehensive study examining how the cure cycle governs phase transitions in a laminate, from viscous to viscoelastic to elastic, and how these transitions influence modulus development, thermal expansion, and cure shrinkage was presented in previous work by Limaye et al. [18_Limaye2024]. These interrelationships were captured using Polymerization–Gelation–Vitrification (PGV) diagrams, which show that polymerization begins when the material reaches its minimum viscosity, gelation occurs at a viscosity of $100\,\text{Pa}\cdot\text{s}$ for the studied resin, and vitrification takes place when the instantaneous glass transition temperature equals the cure temperature. The study further demonstrated that thermal effects depend on both temperature and degree of cure, whereas cure shrinkage is governed solely by resin crosslinking; the coupling between these phenomena is strongest in the interval between gelation and vitrification.
To capture these complex interactions, a computational cure analysis model was developed using ABAQUS and COMPRO in a sequential two-step process encompassing thermo-chemical and stress-deformation analyses. This model provided outputs for temperature, degree of cure, and deformation, and its predictions were verified against existing literature with strong agreement in both degree of cure curves and out-of-plane curvature measurements.
For the simplified flat laminate ($152\,\text{mm}\times 25\,\text{mm}\times 0.3\,\text{mm}$, 2 plies with a $0/90$ layup), the baseline cure cycle employed a steady isothermal dwell at $180^{\circ}\text{C}$ with a ramp-up rate of $2.6^{\circ}\text{C/min}$ and a cool-down rate of $4.846^{\circ}\text{C/min}$, during which cure shrinkage dominated in the visco-elastic phase. In contrast, the Modified Cure Cycle I introduced a non-isothermal dwell with an initial ramp to $140^{\circ}\text{C}$ at $2.6^{\circ}\text{C/min}$, followed by a slower ramp to $180^{\circ}\text{C}$ at $0.315^{\circ}\text{C/min}$ and the same cool-down rate. This modification resulted in thermal expansion that counteracted cure shrinkage during the cooling phase, thereby reducing the overall deformation. Furthermore, two optimization cases, R11 and R21, were evaluated using the NSGA-II genetic algorithm with the goal of minimizing the PID by adapting the slopes of the cure cycle’s two linear ramp segments. In these cases, the heating time was fixed at 1 minute for R11 and 10 minutes for R21, with the constraints of achieving a full cure ($\text{DoC}\geq 0.990$) and ensuring interaction between thermal and shrinkage effects. The optimized cure cycles were found to be 1.001 minutes at $134.36^{\circ}\text{C}$ for R11 and 10.230 minutes at $135.76^{\circ}\text{C}$ for R21, yielding an 8–10% reduction in deformation compared to the baseline cycle.
The cure analysis is carried out as a two-step procedure, consisting of a thermo-chemical step followed by a stress–deformation step. In this framework, a thermo-chemical analysis is first performed to compute the evolving temperature and degree of cure, and these fields are then passed to a subsequent stress–deformation analysis to predict laminate strains and deformation. The key inputs and outputs associated with each stage of this simulation workflow are summarized in [18_Limaye2024]. Cure cycles R11 and R21 were simulated using an initial degree of cure of $\alpha_{i}=0.3$, and the end-of-cycle deformation was compared with experimental measurements (see Table 2 and Figure 4) for validation. The corresponding simulation data are used to assess the performance of the transfer learning approach. Since the deformation at the final time is available experimentally, R11 and R21 serve as validation cases for the FiLM-DeepONet model to evaluate how well transfer learning predicts the deformation history.
@@
(b) Degree of cure and viscosity profiles from the thermo–chemical simulation, together with the corresponding deformation from the stress–deformation simulation, for $\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$.
Figure 6: Simulation data for a laminate subjected to a cure cycle. The top panel combines a parametrized cure temperature profile with an intermediate control point $A=(t_{1},T_{1})$ and fixed start, peak, and end temperatures, and the corresponding imposed temperature histories at different temperature coordinates $T_{c}$ used in simulation data. The bottom panel shows the simulated evolution of degree of cure, viscosity, and deformation for two initial degrees of cure, $\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$.
As shown in Figure 6, a family of simulated temperature profiles is generated by varying an intermediate temperature coordinate $A=(t_{1},T_{1})$ between the fixed start point $(t_{0},T_{\text{start}})=(0.333,20.)$ and peak point $(t_{2},T_{\text{peak}})=(171.658,179.905)$, while keeping the end point $(t_{3},T_{\text{end}})=(205,20.)$ fixed. This intermediate coordinate controls both the duration and the temperature level of the dwell during the heating stage. Figure 6(b) shows the corresponding degree of cure, viscosity, and deformation histories for $\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$ and for different temperature coordinates $T_{c}$, illustrating the influence of the initial degree of cure on the deformation evolution.
@@ 4.1 DeepONet
In this work, DeepONet is configured to learn the mapping from the imposed temperature profile $\mathbf{T}$ of a cure cycle to the degree of cure, viscosity, and deformation history of the laminate. The branch network takes the temperature field $\mathbf{T}$ as input, represented by samples of the cure-cycle temperature profile for different temperature coordinates $T_{c}$ (see Figure 6),
@@ 4.2 FiLM-enhanced DeepONet
As described in Section 4.1, the standard DeepONet architecture learns an operator by combining a branch network that encodes the input function and a trunk network that represents the evaluation coordinates. However, in the present problem, the cure response of the laminate depends not only on the applied temperature history but also on the initial degree of cure $DoC_{0}$. As illustrated in Figure 6(b), variations in $DoC_{0}$ significantly influence the evolution of degree of cure, viscosity, and deformation. To incorporate this additional physical dependence, we augment the branch network with Feature-wise Linear Modulation (FiLM), resulting in a DeepONet architecture. FiLM provides a mechanism for conditioning intermediate branch features on an auxiliary variable—in this case, $DoC_{0}$—allowing the operator representation to adapt dynamically to different initial material states. For a hidden-layer activation $h$ in the branch network, the FiLM transformation is given by
@@
with $\odot$ denoting element-wise multiplication. Through this modulation, the branch network produces a representation of the temperature history that is explicitly conditioned on the initial degree of cure.
@@
where $\mathbf{h}_{d}$, $\mathbf{h}_{v}$, and $\mathbf{h}_{\varepsilon}\in\mathbb{R}^{G}$ encode the operator representations for degree of cure, viscosity, and deformation, respectively. The trunk network takes the query time $t$ as input and produces a shared latent basis:
@@ 5 DeepONet-Based Framework for Modeling Cure-Induced Material Response
Figure 7 presents a schematic of the DeepONet architecture used to predict the degree of cure, viscosity, and deformation of the composite during the cure process. In the first stage, DeepONet is trained on simulation data, the branch network receives the temperature profile, the trunk network takes time as input, and their combination yields the full temporal histories of degree of cure, viscosity, and deformation. In the second stage, the trained DeepONet is employed as a surrogate model for the experiments. The simulated temperature profiles are defined on a common time interval from 0 to $205$ minutes, whereas the experimental temperature profiles span different time windows (e.g., up to approximately $255$, $211.6$, or $227.6$ minutes, depending on the test) and are available only in limited quantity. Training a new network solely on this experimental dataset would be inefficient and prone to overfitting. Instead, a transfer-learning strategy is adopted, the weights learned from the simulation stage are retained, and only the last layer of the branch network is updated so that the predicted deformation matches the experimentally observed final deformation. An additional loss term penalizes the discrepancy between the predicted and measured deformation at the final time, yielding corrected deformation histories for each experimental temperature profile while still leveraging the rich operator learned from the simulations.
@@
As illustrated in Figure 7, the branch network takes as input the discretized temperature field $\mathbf{T}$, sampled at a set of temperature coordinates, together with the initial degree of cure, and outputs three sets of latent coefficients corresponding to the histories of degree of cure, viscosity, and deformation. The trunk network receives the time coordinates and produces a corresponding set of basis functions; the predicted histories are then obtained by combining the branch and trunk outputs. Both branch and trunk networks employ three hidden layers with 20 neurons each and hyperbolic tangent ($\tanh$) activation functions, followed by an output layer with 20 neurons. The model is trained on simulation data using the Adam optimizer with an exponentially decaying learning rate, initialized at $10^{-3}$ and updated according to an ExponentialDecayschedule, for up to $10^{5}$ iterations with early stopping.
[…]
@@
where $\varepsilon=10^{-8}$ is a small constant that ensures numerical stability. The figure also illustrates how FiLM enables the network to clearly distinguish the responses associated with different initial values of the degree of cure.

### v2

@@ 1 Introduction
Along with advances in physics-based modeling, rapid progress in machine learning (ML) has transformed manufacturing research by enabling efficient data-driven modeling, prediction, and optimization [behbahani2022machine, ademujimi2017review, park2025machine, kiyani2023designing]. Deep learning, in particular, has become a central tool across science and engineering. A growing body of work focuses on operator learning, which aims to approximate mappings between input and output functions, a setting that naturally arises in many scientific and engineering applications. For PID in composite structures, a range of ML architectures has been explored. Lin et al. [lin2021use] used neural networks, support vector regression, and $k$-nearest neighbors models to rapidly estimate differences in DoC from process parameters, while Fan et al. [fan2023deep] employed convolutional neural networks to predict PID fields for laminates with varying stacking sequences. Theory-guided and simulation-informed approaches have also been proposed, including theory-guided ML with Gaussian process regression for process optimization [schoenholz2024accelerated], random forests for optimal cure design in co-cured composite structures [2022_Lavaggi], and simulation frameworks that capture competing mechanisms of residual stress evolution [18_Limaye2024]. More recently, Liu et al. [liu2024rapid] combined viscoelastic finite element analysis, feature selection (Random Forest, LASSO), and artificial neural networks for rapid PID prediction, while Zhang et al. [zhang2025deep] introduced deep learning models tailored to complex thermo–chemo–mechanical interactions.
@@
Building on these developments, Feature-wise Linear Modulation (FiLM) [perez2018film] is extended to operator learning. FiLM was originally introduced for visual reasoning and has since been incorporated into architectures such as graph neural networks [brockschmidt2020gnn]. In the present work, a FiLM-augmented DeepONet surrogate model is employed to reconstruct the full histories of DoC, viscosity, and deformation, despite the fact that only the final deformation at the end of the cure cycle is available experimentally. To address the mismatch between the time discretization used in simulations and that available in experiments, a transfer learning strategy [pan2020transfer] is adopted, in which the DeepONet is first trained on high-fidelity simulation data and subsequently fine-tuned using experimental measurements through a loss function that enforces agreement between the predicted and measured final deformation.
@@
While the FiLM-DeepONet framework provides accurate predictions for the evolution of DoC, viscosity, and deformation, performing reliable uncertainty quantification (UQ) for operator learning remains substantially more challenging. The associated inference problem is typically a very high-dimensional Bayesian inverse problem, which renders gradient-based sampling methods prohibitively expensive. Classical Bayesian approaches such as Hamiltonian Monte Carlo (HMC) [neal2012bayesian] and the Laplace approximation [mackay1992practical] can, in principle, deliver high-quality uncertainty estimates, but they scale poorly, even when diagonal approximations of the Hessian are employed. As a practical alternative, Lakshminarayanan et al. [lakshminarayanan2017simple] proposed deep ensembles trained with proper scoring rules, which can yield well-calibrated predictive distributions with only minor modifications to standard training pipelines. However, these ensemble methods have been criticized for lacking a clear Bayesian interpretation [wilson2022evaluating, reich2013nonparametric, pearce2018uncertainty]. To help bridge this gap, Pearce et al. [pearce2020uncertainty] introduced an ensemble scheme based on approximate Bayesian inference, in which model parameters are regularized around samples from a prior-like distribution, leading to more interpretable uncertainty estimates.
@@
Another promising direction is ensemble Kalman inversion (EKI), a gradient-free ensemble method that has been successfully combined with UQ in scientific machine learning [gao2025scalable, pensoneault2025uncertainty]. Owing to its favorable scaling and rapid convergence, EKI is particularly well suited to high-dimensional operator-learning problems. In this work, we therefore employ EKI to train DeepONet while simultaneously obtaining uncertainty estimates. To further reduce the effective parameter dimension and improve computational efficiency, conventional multilayer perceptrons (MLPs) are replaced with Chebyshev enhanced Kolmogorov Arnold Networks (cKANs) within the DeepONet architecture. The cKANs employed here follow a modified DeepOKAN design, in which Chebyshev polynomials are incorporated into Kolmogorov Arnold Networks (KANs) [liu2024kan, shukla2024comprehensive, faroughi2025scientific] within a standard DeepONet framework. Inspired by Kolmogorov networks [sprecher2002space, koppen2002training], KANs are designed to adapt their activation patterns to the input data, thereby enabling more efficient forward evaluations and, consequently, faster generation of ensemble samples. The model developed as part of this study, thus couples operator learning with transfer learning by training a DeepONet surrogate on simulation data and refining it with experimental measurements to recover full deformation fields from limited observations, while EKI-based training simultaneously yields uncertainty estimates, providing experimental predictions that are both accurate and uncertainty-aware.
@@ 2.1 Experimental Approach
The laminate layup was carried out with particular care to ensure consistent ply orientation across all specimens, thereby reducing variability associated with fiber misalignment. Samples fabricated from the selected thermoset prepreg were cured under constant consolidation pressure using one baseline isothermal cycle and two optimized non-isothermal cycles aimed at mitigating PID. The temperature histories of the specimens were recorded in real time and subsequently used to predict the deformation history using a FiLM-enhanced DeepONet.
@@ 2.2 Materials and Methods
White et al. [25_White1993] 1993 AS4 CF/3501-6 Prepreg (42% resin mass) @ 20 °C/min to 350 °C 182.87 (prepreg) 435 (comp. for resin) NA
Kim et al. [21_Kim2002] 2002 AS4 CF/3501-6 Prepreg (37.4% resin mass) @ 20 °C/min to 350 °C 433.7 (comp. for resin) NA
Kim et al. [21_Kim2002] 2002 AS4 CF/3501-6 Prepreg (37.4% resin mass) @ 2 °C/min to 350 °C 456.7 (comp. for resin) NA
Chern et al. [23_Chern2002] 2002 3501-6 Resin @ 20 °C/min $508\pm 19$ 2% cure (manufacturer’s note)
Hargis et al. [20_Hargis2006] 2006 3501-6 Resin @ 20 °C/min $382.5\pm 20$ Significant pre-reaction reported
@@
Based on the values reported in Table 1 and the procedure in Equation (1), $\mathrm{DoC}_{0}$ was estimated to lie within the range of 17.68%–31.21% when the measurement standard deviations were taken into account. Using the reported mean values, $\mathrm{DoC}_{0}$ was calculated to be 24.7%. Accordingly, an initial condition of $\mathrm{DoC}_{0}=0.3$ (30%) was assumed for the simulations. This approximation also accounts for additional curing that may have occurred during the several months between prepreg manufacture and its use in the experiments.
@@ 3 Model and Simulations
A comprehensive study by Limaye et al. [18_Limaye2024] demonstrated that the cure cycle governs phase transitions in polymer laminates, progressing from viscous to viscoelastic and ultimately to elastic behavior. The study further highlighted the critical role of thermo-mechanical interactions in modulus evolution and the development of residual stresses and deformations. These coupled phenomena were represented using Polymerization–Gelation–Vitrification (PGV) plots, which characterize material behavior as a function of the applied cure cycle. To capture these complex interactions, a computational cure analysis framework was developed using ABAQUS and COMPRO in a sequential two-step approach, comprising thermo-chemical analysis followed by stress–deformation analysis. The model predicted temperature evolution, DoC, and laminate deformation. These predictions were validated against existing literature, showing strong agreement in both DoC profiles and deformation measurements.
An optimization scheme was developed and implemented for the R11 and R21 cases using the NSGA-II genetic algorithm, with the objective of minimizing PID. This was accomplished by adjusting the slopes of the cure cycle’s two linear ramp segments, resulting in non-isothermal cure cycles that reduced overall deformation by balancing cure shrinkage with thermal expansion during processing. For these cases, $\mathrm{DoC}_{0}$ was prescribed, while the heating time was fixed at 1 minute for R11 and 10 minutes for R21. The optimization was performed subject to the constraints of achieving a high final cure state ($\mathrm{DoC}\geq 0.990$) and promoting interaction between thermal expansion and chemical shrinkage effects.
In the present work, the cure model is regenerated to evaluate non-isothermal cure cycles while accounting for material pre-cure through the prescribed value of $\mathrm{DoC}_{0}$. The end-of-cycle deformation predicted for the simulated cure cycles R11 and R21 is compared against experimental measurements (see Table 2 and Figure 4) to validate the model. The corresponding simulation data are subsequently used to assess the performance of the transfer learning approach. Because the final-time deformation is available experimentally, the R11 and R21 cases serve as validation benchmarks for the FiLM-DeepONet model to evaluate the accuracy of deformation-history predictions enabled by transfer learning.
As shown in Figure 6(a), a family of simulated temperature profiles is generated by varying an intermediate temperature coordinate $A=(t_{1},T_{1})$ between the fixed start point $(t_{0},T_{\mathrm{start}})=(0.333,20)$ and peak point $(t_{2},T_{\mathrm{peak}})=(171.658,179.905)$, while keeping the end point $(t_{3},T_{\mathrm{end}})=(205,20)$ fixed. This intermediate coordinate controls both the duration and the temperature level of the dwell during the heating stage. Figure 6(b) shows the corresponding histories of DoC, viscosity, and deformation for different choices of the temperature coordinate $A$, illustrating the influence of $\mathrm{DoC}_{0}$ on the deformation evolution.
@@
(b) DoC and viscosity profiles from the thermo–chemical simulation, together with the corresponding deformation from the stress–deformation simulation, for $\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$.
Figure 6: Simulation data for a laminate subjected to a cure cycle. The top panel combines a parametrized cure temperature profile with an intermediate control point $A=(t_{1},T_{1})$ and fixed start, peak, and end temperatures, and the corresponding imposed temperature histories at different temperature coordinates $T_{c}$ used in simulation data. The bottom panel shows the simulated evolution of DoC, viscosity, and deformation for two initial degrees of cure, $\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$.
@@ 4.1 DeepONet
In this work, DeepONet is configured to learn the mapping from the imposed temperature profile $\mathbf{T}$ of a cure cycle to the DoC, viscosity, and deformation history of the laminate. The branch network takes the temperature field $\mathbf{T}$ as input, represented by samples of the cure-cycle temperature profile for different temperature coordinates $T_{c}$ (see Figure 6(a)),
@@ 4.2 FiLM-enhanced DeepONet
As described in Section 4.1, the standard DeepONet architecture learns an operator by combining a branch network, which encodes the input function, with a trunk network, which represents the evaluation coordinates. However, in the present problem, the laminate response depends not only on the applied temperature history but also on $\mathrm{DoC}_{0}$. As illustrated in Figure 6(b), variations in $\mathrm{DoC}_{0}$ significantly influence the evolution of $\mathrm{DoC}$, viscosity, and deformation.
To incorporate this additional physical dependence, we augment the branch network with Feature-wise Linear Modulation (FiLM), resulting in a DeepONet architecture. FiLM provides a mechanism for conditioning intermediate branch features on an auxiliary variable—in this case, $DoC_{0}$—allowing the operator representation to adapt dynamically to different initial material states. For a hidden-layer activation $h$ in the branch network, the FiLM transformation is given by
@@
with $\odot$ denoting element-wise multiplication. Through this modulation, the branch network produces a representation of the temperature history that is explicitly conditioned on $\mathrm{DoC}_{0}$.
@@
where $\mathbf{h}_{d}$, $\mathbf{h}_{v}$, and $\mathbf{h}_{\varepsilon}\in\mathbb{R}^{G}$ encode the operator representations for DoC, viscosity, and deformation, respectively. The trunk network takes the query time $t$ as input and produces a shared latent basis:
@@ 5 DeepONet-Based Framework for Modeling Cure-Induced Material Response
Figure 7 presents a schematic of the DeepONet architecture used to predict the DoC, viscosity, and deformation of the composite during the cure process. In the first stage, DeepONet is trained on simulation data, the branch network receives the temperature profile, the trunk network takes time as input, and their combination yields the full temporal histories of DoC, viscosity, and deformation. In the second stage, the trained DeepONet is employed as a surrogate model for the experiments. The simulated temperature profiles are defined on a common time interval from 0 to $205$ minutes, whereas the experimental temperature profiles span different time windows (e.g., up to approximately $255$, $211.6$, or $227.6$ minutes, depending on the test) and are available only in limited quantity. Training a new network solely on this experimental dataset would be inefficient and prone to overfitting. Instead, a transfer-learning strategy is adopted, the weights learned from the simulation stage are retained, and only the last layer of the branch network is updated so that the predicted deformation matches the experimentally observed final deformation. An additional loss term penalizes the discrepancy between the predicted and measured deformation at the final time, yielding corrected deformation histories for each experimental temperature profile while still leveraging the rich operator learned from the simulations.
@@
As illustrated in Figure 7, the branch network takes as input the discretized temperature field $\mathbf{T}$, sampled at a set of temperature coordinates, together with $\mathrm{DoC}_{0}$, and outputs three sets of latent coefficients corresponding to the histories of DoC, viscosity, and deformation. The trunk network receives the time coordinates and produces a corresponding set of basis functions; the predicted histories are then obtained by combining the branch and trunk outputs. Both branch and trunk networks employ three hidden layers with 20 neurons each and hyperbolic tangent ($\tanh$) activation functions, followed by an output layer with 20 neurons. The model is trained on simulation data using the Adam optimizer with an exponentially decaying learning rate, initialized at $10^{-3}$ and updated according to an ExponentialDecayschedule, for up to $10^{5}$ iterations with early stopping.
Figure 8: Comparison of FiLM-DeepONet predictions with true simulation results at three temperature coordinates, including DoC for two initial conditions ($\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$), viscosity, and deformation. $\mathrm{DoC}_{0}$ is provided as a conditioning input to the branch network.
Figure 8 compares DeepONet predictions with reference simulation results for $\mathrm{DoC}$, viscosity, and deformation over time under four temperature profiles, using two initial conditions ($\mathrm{DoC}_{0}=0.001$ and $\mathrm{DoC}_{0}=0.3$). In the FiLM-enhanced architecture, $\mathrm{DoC}_{0}$ conditions the branch network, enabling the model to capture how the coupled thermochemical and mechanical response varies with the initial cure state. The close agreement between predictions and reference data demonstrates both the accuracy and generalization capability of the trained DeepONet. The rightmost panel further indicates that $\mathrm{DoC}_{0}$ strongly influences the final deformation pattern, justifying its inclusion as a conditioning variable. As shown in the figure, viscosity grows approximately exponentially with time, leading to steep gradients and a wide dynamic range. Such highly skewed data can be challenging for neural networks to learn directly. To mitigate this, viscosity was log-transformed prior to training,
@@
where $\varepsilon=10^{-8}$ is a small constant introduced for numerical stability. The figure also illustrates that FiLM enables the network to clearly distinguish responses corresponding to different initial degrees of cure.